# CAUS: A Dataset for Question Generation based on Human Cognition Leveraging Large Language Models

### v1

@@ Discussion
In this study, we designed a novel dataset based on human cognition along with a pipeline generating human-like questions and question classification structure. Our approach leveraged cutting-edge tools such as GPT-4 model, and we performed in-depth evaluations to assess its performance. We made the controlled uncertainty with scene descriptions that offer a text version of Out Of Distribution (OOD). The uncertainty was fine-tuned to remain predictable, though not as tightly structured as in a game context. The main focus was understanding the LLM’s capability to identify and inquire about uncertain elements within a given context. We also explored whether LLM can capture the nuanced content and format of given questions by the question classification.
To wrap up, we revisit the two questions posed at the outset. Firstly, we asked: Are LLMs inherently incompetent in asking questions? Contrary to this assumption, we found that LLMs show the ability to generate appropriate questions. Through our endeavors, we have determined that it is indeed possible to extract a suitable set of questions from LLMs by offering proper context and instruction. Moreover, LLM’s high question classification performance also revealed that LLM is good at predicting the properties of questions as if the model understands the questioner’s purpose and motivation. In conclusion, LLMs seem to have the potential to ask and grasp the nuance.
@@ Acknowledgments
Anonymous CogSci submission. This will be updated in the final version.

### v2

@@ Discussion
In this study, we designed a novel dataset based on human cognition, along with a pipeline generating human-like questions and question classification structures. Our approach leveraged cutting-edge tools such as the GPT-4 model, and we performed in-depth evaluations to assess its performance. We made the controlled uncertainty with scene descriptions that offer a text version of Out Of Distribution (OOD). The uncertainty was fine-tuned to remain predictable, though not as tightly structured as in a game context. The main focus was understanding the LLM’s capability to identify and inquire about uncertain elements within a given context. We also explored whether LLM can capture the nuanced content and format of given questions by the question classification.
To wrap up, we revisit the two questions posed at the outset. Firstly, we asked: Are LLMs inherently incompetent at asking questions? Our findings challenge this prevailing idea, demonstrating that LLMs can generate appropriate questions with proper context and instruction. Moreover, LLM’s high question classification performance also revealed that LLM is good at predicting the properties of questions, as if the model understands the questioner’s purpose and motivation. In conclusion, LLMs seem to have the potential to ask and grasp the nuance.
@@ Acknowledgments
We deeply thank the reviewers for providing kind and helpful comments. And this work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. 20220-00951, Development of Uncertainty-Aware Agents Learning by Asking Questions)